Two hundred trolls divided a block of gold evenly. What percent of
the block of gold did each troll receive?
a) 200% b) 1/200% c) c) 2% d) 0.5% e) 0.2%

Answer by stanbon(57387)

Two hundred trolls divided a block of gold evenly. What percent of
the block of gold did each troll receive?
------------
Percent means per hundred.
------------
1/200 = (1/2)/100 = (1/2)percent = 0.5%
============================================
Cheers,
Stan H.
a) 200% b) 1/200% c) c) 2% d) 0.5% e) 0.2%